OpenStudy (anonymous):

Can someone help me out? 2. Evaluate the expression for a = –1 and b = 5. Show your work. 5a – 7b + b^2 This is how i solved it i'm just seeing if i did it right or if i messed up somewhere throughout the problem 5+ -1 -7(5) + (5)^2 -35 25 4+ -35 25 4+ -35 = -31 -31+ 25 = -6

OpenStudy (abdullahm):

You messed up.

OpenStudy (abdullahm):

\(\sf\Large 5a – 7b + b^2\) where \(\sf\Large a=-1 \\\sf\Large b=5\)

OpenStudy (abdullahm):

So simply just plug in the numbers: \(\sf\Large 5(-1) -7(5)+(5)^2=?\)

OpenStudy (abdullahm):

Remember, it's multiplication we are doing. Not addition or subtraction.

OpenStudy (anonymous):

Alright is this correct 5+( -1) -7(5) + (5)^2 -35 25 - 5 -35 25 -5+ -35 = 40 -40+ 25 = 15

OpenStudy (abdullahm):

almost correct, but \(\sf\Large -40+25 = \color{red}{-}15\) because \(\sf\Large -40+25 = 25-40\)

OpenStudy (anonymous):

5+( -1) -7(5) + (5)^2 -35 25 - 5 -35 25 -5+ -35 = -40 25 + -40 = 15 How about now?

OpenStudy (abdullahm):

you have all the steps correct except the last one \(\sf\Large 25+ (-40) = -15\) \(\sf\Large -40 + 25 = -15\) It's \(\bf \color{red}{-}\)15 It's a negative 15...

OpenStudy (anonymous):

Oh didn't notice that >~< I'm sorry.

OpenStudy (abdullahm):

It's ok, we should learn from our mistakes. :) And Welcome to OpenStudy! @Ashtaroth

OpenStudy (anonymous):

Thanks. I think this place is going to help me big time with my algebra.

OpenStudy (anonymous):

Could you help me with another?

OpenStudy (abdullahm):

sure close this question and post another question :)